OpenStudy (dmezzullo):

Consider the line that passes through the points (0, 4) and (-1, 3). Below are two different equations. Identify the true statement concerning both equations. Equation #1 Equation #2 y - 4 = 1(x - 0) y - 3 = 1(x - 1) Only equation #1 represents the line that passes through the two given points. Only equation #2 represents the line that passes through the two given points. Both equation #1 and equation #2 represent the line that passes through the two given points. Neither equation #1 nor equation #2 represents the line that passes through the two given points.

OpenStudy (dmezzullo):

@ceb105

OpenStudy (anonymous):

Could you rewrite the two equations more clearly seperated so I can see the end of the first and the beginning of the second?

OpenStudy (dmezzullo):

Equation #1 y - 4 = 1(x - 0)

OpenStudy (dmezzullo):

Equation #2 y - 4 = 1(x - 0)

OpenStudy (anonymous):

Okay so this question is solely involving substituting in your values of (x,y) and seeing if the equation is satisfied, if an equation satifies both values of (-1,3) and (0,4) then it is representing a line that passes through those points.

OpenStudy (dmezzullo):

So then it would be c?

OpenStudy (anonymous):

Try substituting in your values into the #2 equation, I think you may have made some errors

OpenStudy (anonymous):

The answer is only #1 represents the line passing through both points

OpenStudy (dmezzullo):

OHHH OK